# arXiv:2603.00942v1[math.DG] 01 Mar 2026

### v1

@@ 6 Appendix B
Assume that the potential function $\phi$ has Property 1. The eigenfunction corresponding to the first Dirichlet eigenvalue of the Witten-Laplacian on the geodesic ball $\mathscr{B}_{n}(q^{-},r_{0})$ (resp., $\mathscr{B}_{n}(q^{+},r_{0})$) may be chosen to be nonnegative and is a radial function $\psi(s)$ satisfying $\psi^{\prime}(s)<0$ for $0<s\leq r_{0}<\min\{\ell(q),l\}$, with $\ell(q)$ defined as in (2.7) (resp., for $0<s\leq r_{0}<\min\{\mathrm{inj}(q),l\}$).
@@
where $\Delta_{\mathbb{S}^{n-1}}$ denotes the Laplacian on $\mathbb{S}^{n-1}$ with respect to the round metric. Using a similar argument to Courant’s nodal domain theorem for the Laplacian shown in [15, Chapter I], one knows that the first Dirichlet eigenvalue $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ of the Witten-Laplacian is simple (i.e. the eigenspace of this eigenvalue has dimension $1$) and its eigenfunction $\psi$ would not change sign on $\mathscr{B}_{n}(q^{-},r_{0})$. One can also check the proof of our recent work [22, Theorem 1.2] for details if he or she is not familiar with this content. Without loss of generality, one may choose that $\psi$ is nonnegative on $\mathscr{B}_{n}(q^{-},r_{0})$. More precisely, $\psi$ is strictly positive within $\mathscr{B}_{n}(q^{-},r_{0})$ and vanishes on the boundary $\partial\mathscr{B}_{n}(q^{-},r_{0})$. Since the dimension of the eigenspace corresponding to $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ is $1$, it follows that $\psi$ is radial and satisfies
@@
Now, together with Lemma 6.1, we wish to prove Theorem 1.17 by using a similar argument to that in the proof of [31, Theorem 3.6].
@@
$\frac{1}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]$ $=$ $-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)$ (6.5)
$\qquad+(n-1)\frac{f(s)}{J(s,\xi)}\left[\frac{J(s,\xi)}{f(s)}\right]^{\prime}\frac{d\psi}{ds}$
$\geq$ $-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)\qquad$
by using the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6. Substituting (6.5) into (6) yields
@@
Let $\Psi$ be an eigenfunction corresponding to $\lambda_{1,\phi}(\Omega)$, with $\Psi|_{\Omega}>0$ and $\Psi|_{\partial\Omega}=0$. This choice of the eigenfunction $\Psi$ can be assured by [22, Theorem 1.2], i.e. the number of nodal domains of the eigenfunction $\Psi$ is $1$, which implies $\Psi$ would not change sign on $\Omega$. Set
@@
We shall use a similar argument to that in the proof of [31, Theorem 4.4]. Let $\psi(s):[0,r_{0}]\rightarrow[0,\infty)$ be a nonnegative radial eigenfunction of $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$ of $M^{+}$. By Lemma 6.1, one knows that the eigenfunction $\psi$ satisfies (6.1), with $\psi^{\prime}(0)=\psi(r_{0})=0$, $\psi\geq 0$ on $0,r_{0})$, and $\frac{d\psi}{ds}<0$ on $(0,r_{0})$. Define a function $\mathcal{F}:\overline{B(q,r_{0})}\rightarrow[0,\infty)$ by
@@
for $(s,\xi)\in[0,r_{0})\times S_{q}^{n-1}$. Then by a straightforward calculation and using the characterization (1.13), one can get
@@
Assume that the potential function $\phi$ has Property 1. The eigenfunction corresponding to the first Dirichlet eigenvalue of the weighted $p$-Laplacian ($1<p<\infty$) on the geodesic ball $\mathscr{B}_{n}(q^{-},r_{0})$ may be chosen to be nonnegative and is a radial function $\psi(s)$ satisfying $\psi^{\prime}(s)<0$ for $0<s\leq r_{0}<\min\{\ell(q),l\}$, with $\ell(q)$ defined as in (2.7).
@@
$(p-1)|\psi^{\prime}(s)|^{p-2}\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)\right]|\psi^{\prime}(s)|^{p-2}\psi^{\prime}(s)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\lambda^{\phi}_{1,p}(\mathscr{B}_{n}(q^{-},r_{0}))|\psi(s)|^{p-2}\psi(s)=0$ (6.8)
and the boundary condition $\psi^{\prime}(0)=0$ and $\psi(r_{0})=0$. Set $\alpha(s)=e^{\int_{\epsilon}^{s}\left[(n-1)\frac{f^{\prime}(t)}{f(t)}-\phi^{\prime}(t)\right]dt}$, $\epsilon\rightarrow 0^{+}$, and then it is not hard to know that (6) becomes
@@
We wish to use a similar argument to that in the proof of [47, Theorem 3.2]. Let $\psi$ be an eigenfunction of the Dirichlet eigenvalue $\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$, which, by Lemma 6.5, is simple, radial and can be chosen to be nonnegative. Since $\psi\circ r$ vanishes on the boundary $\partial B(q,r_{0})$, where $r(x)=d_{M^{n}}(q,x)$ measures the radial distance from $q$ to any point $x\in\partial B(q,r_{0})$, it is not hard to know that $\psi\circ r\in W^{1,p}_{0,\phi}(B(q,r_{0}))$ and
@@
$\int\limits_{0}^{a(\xi)}|\psi^{\prime}(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds=-\psi(s)|\psi^{\prime}(s)|^{p-1}e^{-\phi}f^{n-1}(s)\theta(s,\xi)\Big|_{0}^{a(\xi)}+$
$\int\limits_{0}^{a(\xi)}\frac{\psi(s)}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[|\psi^{\prime}(s)|^{p-1}f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\qquad\quad$ (6.11)
@@
Together with (6), one has from (6) that
@@
by using the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6. Substituting the above inequality into (6) results in
@@
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$. This is exactly the assertions in [47, Theorem 3.2]. If furthermore the curvature assumption was strengthened to be $\mathrm{Ric}(M^{n})\geq(n-1)K$ for some constant $K$, then the eigenvalue comparison (6.14) would reduce to
@@
This spectral asymptotical property on the unit Euclidean (n1)(n-1)-sphere $\mathbb{S}^{n-1}$ has been extended to the first Dirichlet eigenvalue of the Laplacian on “closed” spherically symmetric manifolds (see [31, lemma2.5]), and later to the first Dirichlet eigenvalue of the $p$-Laplacian ($1<p<\infty$) on “closed” spherically symmetric manifolds (see [47, Lemma 2.2]). It is natural to ask whether or not these extensions can be further improved to the case of weighted $p$-Laplacian. In fact, by using the same cut-off continuous functions as those two in the proof of [47, Lemma 2.2], one has:
@@
The reason why the same cut-off continuous functions still work in the weighted case is that $e^{-\phi}$ is always bounded on the geodesic ball $\mathscr{B}_{n}(q^{\ast},r)\subseteq M^{\ast}$, which leads to the result that all integrals appeared in the proof of [47, Lemma 2.2] (i.e. corresponding to the non-weighted case) have no essential difference with those would appear in the weighted case.

### v2

@@ 6 Appendix B
Assume that the potential function $\phi$ has Property 1. The eigenfunction corresponding to the first Dirichlet eigenvalue of the Witten-Laplacian on the geodesic ball $\mathscr{B}_{n}(q^{-},r_{0})$ (resp., $\mathscr{B}_{n}(q^{+},r_{0})$) may be chosen to be nonnegative and is a radial function $\psi(s)$ satisfying $\psi^{\prime}(s)<0$ for $0<s\leq r_{0}<\min\{\ell(q),l\}$, with $\ell(q)$ defined as in (2.7) (resp., for $0<s\leq r_{0}<\min\{\mathrm{inj}(q),l\}$, with $\mathrm{inj}(q)$ defined by (2.5)).
@@
where $\Delta_{\mathbb{S}^{n-1}}$ denotes the Laplacian on $\mathbb{S}^{n-1}$ with respect to the round metric. Using a similar argument to Courant’s nodal domain theorem for the Laplacian shown in [16, Chapter I], one knows that the first Dirichlet eigenvalue $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ of the Witten-Laplacian is simple (i.e. the eigenspace of this eigenvalue has dimension $1$) and its eigenfunction $\psi$ would not change sign on $\mathscr{B}_{n}(q^{-},r_{0})$. One can also check the proof of our recent work [23, Theorem 1.2] for details if he or she is not familiar with this content. Without loss of generality, one may choose that $\psi$ is nonnegative on $\mathscr{B}_{n}(q^{-},r_{0})$. More precisely, $\psi$ is strictly positive within $\mathscr{B}_{n}(q^{-},r_{0})$ and vanishes on the boundary $\partial\mathscr{B}_{n}(q^{-},r_{0})$. Since the dimension of the eigenspace corresponding to $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ is $1$, it follows that $\psi$ is radial and satisfies
@@
Now, together with Lemma 6.1, we wish to prove Theorem 1.17 by using a similar argument to that in the proof of [33, Theorem 3.6].
@@
$\qquad\qquad\qquad\qquad\qquad=-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(n-1)\frac{f(s)}{J(s,\xi)}\left[\frac{J(s,\xi)}{f(s)}\right]^{\prime}\frac{d\psi}{ds}$
$\qquad\qquad\qquad\qquad\qquad\geq-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)\qquad$ (6.5)
by using the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6. Substituting (6) into (6) yields
@@
Let $\Psi$ be an eigenfunction corresponding to $\lambda_{1,\phi}(\Omega)$, with $\Psi|_{\Omega}>0$ and $\Psi|_{\partial\Omega}=0$. This choice of the eigenfunction $\Psi$ can be assured by [23, Theorem 1.2], i.e. the number of nodal domains of the eigenfunction $\Psi$ is $1$, which implies $\Psi$ would not change sign on $\Omega$. Set
@@
We shall use a similar argument to that in the proof of [33, Theorem 4.4]. Let $\psi(s):[0,r_{0}]\rightarrow[0,\infty)$ be a nonnegative radial eigenfunction of $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$ of $M^{+}$. By Lemma 6.1, one knows that the eigenfunction $\psi$ satisfies (6.1), with $\psi^{\prime}(0)=\psi(r_{0})=0$, $\psi\geq 0$ on $0,r_{0})$, and $\frac{d\psi}{ds}<0$ on $(0,r_{0})$. Define a function $\mathcal{F}:\overline{B(q,r_{0})}\rightarrow[0,\infty)$ by
@@
for $(s,\xi)\in[0,r_{0})\times S_{q}^{n-1}$. Then by a straightforward calculation in geodesic spherical coordinates around $q$ and using the characterization (1.13), one can get
@@
Assume that the potential function $\phi$ has Property 1. The eigenfunction corresponding to the first Dirichlet eigenvalue of the weighted $p$-Laplacian ($1<p<\infty$) on the geodesic ball $\mathscr{B}_{n}(q^{-},r_{0})$ may be chosen to be nonnegative and is a radial function $\psi(s)$ satisfying $\psi^{\prime}(s)<0$ for $0<s\leq r_{0}<\min\{\ell(q),l\}$, with $\ell(q)$ defined as in (2.7). The conclusion is still valid for the eigenfunction belonging to the first Dirichlet eigenvalue of the weighted $p$-Laplacian on the geodesic ball $\mathscr{B}_{n}(q^{+},r_{0})$ for $0<s\leq r_{0}<\min\{\mathrm{inj}(q),l\}$, with $\mathrm{inj}(q)$ defined by (2.5).
@@
$(p-1)|\psi^{\prime}(s)|^{p-2}\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)\right]|\psi^{\prime}(s)|^{p-2}\psi^{\prime}(s)$ (6.8)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\lambda^{\phi}_{1,p}(\mathscr{B}_{n}(q^{-},r_{0}))|\psi(s)|^{p-2}\psi(s)=0$
and the boundary condition $\psi^{\prime}(0)=0$ and $\psi(r_{0})=0$. Set $\alpha(s)=e^{\int_{\epsilon}^{s}\left[(n-1)\frac{f^{\prime}(t)}{f(t)}-\phi^{\prime}(t)\right]dt}$, $\epsilon\rightarrow 0^{+}$, and then it is not hard to know that (6.8) becomes
@@
We wish to use a similar argument to that in the proof of [49, Theorem 3.2]. Let $\psi$ be an eigenfunction of the Dirichlet eigenvalue $\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$, which, by Lemma 6.5, is simple, radial and can be chosen to be nonnegative. Since $\psi\circ r$ vanishes on the boundary $\partial B(q,r_{0})$, where $r(x)=d_{M^{n}}(q,x)$ measures the radial distance from $q$ to any point $x\in\partial B(q,r_{0})$, it is not hard to know that $\psi\circ r\in W^{1,p}_{0,\phi}(B(q,r_{0}))$ and
@@
$\int\limits_{0}^{a(\xi)}|\psi^{\prime}(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds=-\psi(s)|\psi^{\prime}(s)|^{p-1}e^{-\phi}f^{n-1}(s)\theta(s,\xi)\Big|_{0}^{a(\xi)}+$ (6.11)
$\int\limits_{0}^{a(\xi)}\frac{\psi(s)}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[|\psi^{\prime}(s)|^{p-1}f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\qquad\quad$
@@
Together with (6.8), one has from (6) that
@@
by using the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6. Substituting the above inequality into (6.11) results in
@@
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$. This is exactly the assertions in [49, Theorem 3.2]. If furthermore the curvature assumption was strengthened to be $\mathrm{Ric}(M^{n})\geq(n-1)K$ for some constant $K$, then the eigenvalue comparison (6.14) would reduce to
@@
This spectral asymptotical property on the unit Euclidean (n1)(n-1)-sphere $\mathbb{S}^{n-1}$ has been extended to the first Dirichlet eigenvalue of the Laplacian on “closed” spherically symmetric manifolds (see [33, lemma2.5]), and later to the first Dirichlet eigenvalue of the $p$-Laplacian ($1<p<\infty$) on “closed” spherically symmetric manifolds (see [49, Lemma 2.2]). It is natural to ask whether or not these extensions can be further improved to the case of weighted $p$-Laplacian. In fact, by using the same cut-off continuous functions as those two in the proof of [49, Lemma 2.2], one has:
@@
The reason why the same cut-off continuous functions still work in the weighted case is that $e^{-\phi}$ is always bounded on the geodesic ball $\mathscr{B}_{n}(q^{\ast},r)\subseteq M^{\ast}$, which leads to the result that all integrals appeared in the proof of [49, Lemma 2.2] (i.e. corresponding to the non-weighted case) have no essential difference with those would appear in the weighted case.